# The connection of polaritonic chemistry with the physics of a spin glass

### v1

@@ 4.2.4 Polaritonic resonance picture revised
As we know from a set of different experiments addressing a variety of chemical reactions in cavities (see e.g. Sec. 3.2), resonance effects play a major role when modifying chemical properties under collective VSC. 22, 103, 104 Thus it comes as no surprise that periodic feedback effects between the frustrated (off-equilibrium) electronic structure, the nuclear and cavity degrees of freedom and the thermal bath re-appear in a holistic theoretical description. 5 While considerable theoretical efforts went into investigating resonance effects with polaritonic reaction-rate theories 67, 105, 106, 107, 108, 109, 110 or few-molecule ab-initio simulations 111, 112, 113, we believe that the polarization-glass concept provides a necessary link between the macroscopic (rates) and microscopic (few-molecule) chemical pictures. It opens up the door to identify prerequisites to observe chemical changes in a cavity from an ab-initio perspective. In more detail, we anticipate a scale-connecting resonance mechanism under collective VSC of a form similar to the following: Assuming that the losses of the cavity are sufficiently small (an effective cavity mode is distinguished), a dynamic instability (phase transition) of the dressed electronic structure can emerge for sufficiently strong collective couplings. The resulting polarization-glass phase introduces dynamic correlations between the electronic structures of the molecules. The long-lived correlations imply modifications of the standard (thermal) fluctuation-dissipation relations for the dressed electronic structure, which implies a different out-of-equilibrium description. Eventually, the dynamic (memory-dependent) electronic correlations will act back on the dynamics of the photon displacement field and the nuclei. In that regard, the dynamically frustrated nature of the electronic structure bridges not only length and time-scales, but it also breaks the conservative nature (in a system subspace) of the classical forces of Eq. (5). While the magnitude of the cavity-induced polarizations might be very small, their correlated and long-lived nature may still be sufficient to introduce stochastic resonance phenomena 5 at least for certain chemical systems and thus modifies their respective chemical properties. Clearly, the ab-initio picture of cavity-induced resonance effects remains vague at the moment and considerable theoretical and experimental effort will be required to unravel it and make it more quantitative and predictive. However, we believe that the spin-glass-like nature of the dressed electronic structure provides the most realistic and plausible theoretical framework that sets the necessary seed (instability mechanism) to trigger the resonance effects that have been observed in polaritonic experiments. Therefore, we consider the connection between polaritonic chemistry and the physics of a spin glass an excellent starting point to not only better understand current experimental findings, but also to stimulate novel theoretical and experimental directions. This should allow us to reach a holistic understanding of strongly-coupled light and matter systems. Moreover, our connection provides an interesting new perspective to the applicability of the field of spin glasses beyond condensed matter systems.
@@
(a) Connecting equilibrium and off-equilibrium picture for a spin (black) and polarization (light blue) glass. Left: Equilibrium order parameter distribution of a SK model (black) taken from Ref. 114 in comparison with deviations of $q_{EA}(t)$ in a cavity. Right: In spin glasses two different (weak) off-equilibrium regimes (stationary and aging) can be distinguished for a small external field perturbation $h^{\prime}$, applied after a waiting time $t_{w}$. The stationary regime is governed by a linear fluctuation dissipation relation, which terminates at $C=q_{EA}$ given by the linear response susceptibility $S(q_{EA})=T\chi_{LR}$. In contrast, the aging regime is governed by the modified fluctuation dissipation relations in Eq. (49). It is bounded by the full thermal equilibrium susceptibility $S(0)=T\chi_{eq}$ at $C=0$. Notice that $\chi_{eq}\neq\chi_{LR}$ indicates replica symmetry breaking, whereas $\chi_{eq}=\chi_{LR}$ corresponds to normal hysteresis effects that do not require the complex theory of spin glasses to describe its equilibrium properties (dashed green). The slope of the curve can be interpreted as an effective temperature increase (see Eq. (51)) (red). The dynamic feedback effects in a cavity suggest the breakdown of the linear fluctuation-dissipation relations already within the stationary correlation regime (light-blue).
@@
(b) Experimental off-equilibrium measurement: The breakdown of fluctuation-dissipation relations (aging) measured the after cooling CdCr1.7In0.3S4 below the critical spin glass temperature of $T_{g}=16.2$ K. Bold symbols indicate the measured relaxation-correlation curve for different (finite) waiting times $t_{w}\in\{100,200,500,1000,2000\}$ [s]. Open symbols show the extrapolation to infinite waiting times $t_{w}\rightarrow\infty$The illustration was modified based on Fig. 7 Ref. 75, which contains experimental data from Ref. 102.
@@ 5 Summary and Conclusion
In this work we have established a novel connection between two separated and disconnected research areas: The emergent field of polaritonic chemistry and the long and widely studied complex field of spin glasses and rare events in statistical mechanics. The mapping of the static electronic structure problem of a dilute ensemble under collective VSC, onto the spin-glass model of Sherrington and Kirkpatrick suggests the emergence of a cavity-induced phase transition. The static mapping onto a spin glass indicates the following for the dressed electronic-structure problem:
@@
    Polarization ordering (polarization-glass phase transition) emerges for sufficiently strong collective light-matter couplings to the cavity modes.
@@
    Thermal and non-adiabatic effects can start to play a role for the collective electronic structure problem in a cavity, even though these effects can be usually disregarded for an ensemble in free-space.
@@
    Within the (static) polarization-glass phase, extremely long correlation times and frustration effects emerge, that do not only change equilibrium properties, but also give rise to off-equilibrium aging effects mediated by the long-range cavity modes, which can be re-interpreted as an effective heating of the system.
Overall the standard SK model of a spin glass suggests an instability of the dressed electronic subsystem altering the established temporal and spatial scales to understand chemistry on a molecular scale. Eventually, relaxing the restrictive static two-level approximation in the SK mapping breaks with the established picture of spin glasses. The non-perturbative feedback effect between the dressed electronic problem and the oscillating nuclei and cavity field introduces dynamically frustrated correlations, which require novel concepts and theoretical methods. For instance, the non-monotonic behavior of the correlations in the dynamical situation (see Fig. 9) modifies the simple heating-picture for off-equilibrium effects and could allow cooling as well. Such effects have recently been observed experimentally for collective strong-coupling situations 115. The numerical results suggest the following dynamical features of a polarization glass at the onset of the instability (adiabatic electronic ground-state approximation):
@@
    For a fixed time-step, the polarizations are anti-aligned and narrowly distributed, which gives rise to significant dynamical frustration/correlation effects with zero overall polarization (absence of chemical shifts in NMR experiments 27).
@@
    The dynamical feedback between cavity, nuclei and electronic structure introduces a configuration-dependent (time-dependent) Edwards-Anderson order parameter.
@@
    The cavity-induced polarizations oscillate and are correlated over very long time-scales, which modifies the thermal fluctuation-dissipation relations. However, in the adiabatic limit no waiting time dependency has been detected with our present simulation capabilities and thus aging effects have not (yet) been verified under VSC.
Simulations beyond the adiabatic approximation would be required to fully enter the polarization glass phase, which will become computationally extremely demanding (solving the time-dependent Schrödinger equation and including temperature). However, from the SK model, one can anticipate that spontaneous replica symmetry breaking may also start to play a role. Therefore, one would expect the emergence of dynamical aging effects and thus a cavity-induced time-reversal symmetry breaking. This complicates the theoretical picture even further, particularly if reaching for an analytic understanding in the large-$N$ limit. However, cavity-induced time-reversal symmetry breaking could also open novel pathways for the sub-field of chiral polaritonics. 116 There the common aim is to reach cavity-induced enantioselctivity by explicitly parity-violating cavity polarizations. 117, 112, 118, 119, 120, 121, 122, 123 However, from a fundamental theoretical aspect similar effects may also be reached by spontaneous symmetry breaking instead. 124, 125 Eventually, from a theoretical chemistry perspective, the even more pressing theoretical question is how the feedback between the electronic, nuclear and cavity subsystems can give rise to the experimentally observed stochastic resonance phenomena, as suggested in Sec. 4.2.4. We expect that the instability and thus the dynamical frustration of the dressed electronic structure acts as as seed such that stochastic resonances emerge, i.e., a synchronization of the nuclear/cavity dynamics builds up. However, this picture remains to be proved/disproved. Yet if validated (or supplemented) it could be used as a guiding principle for the design of experimental VSC setups and for the development of effective models of polaritonic chemistry.
@@
We gratefully acknowledge all members of our ERC synergy grant proposal team (Tal Schwartz, Thomas Ebbesen, Abraham Nitzan, Sharly Fleischer, Cyriaque Genet and Maxim Sukharev) for inspiring discussions and helpful comments. This work was made possible through the support of the RouTe Project (13N14839), financed by the Federal Ministry of Education and Research (Bundesministerium für Bildung und Forschung (BMBF)) and supported by the European Research Council (ERC-2015-AdG694097), the Cluster of Excellence “CUI: Advanced Imaging of Matter” of the Deutsche Forschungsgemeinschaft (DFG), EXC 2056, project ID 390715994 and the Grupos Consolidados (IT1249-19). The Flatiron Institute is a division of the Simons Foundation.).
{suppinfo}

### v2

@@ 5.2 Non-canonical nuclear dynamics / stochastic resonances
In that regard, the spin glass nature of the electronic-structure bridges not only different length and time-scales, but it also breaks the conservative nature of the classical forces of Eq. (5). While the magnitude of the cavity-induced non-equilibrium aging dynamics might be very small, their correlated and long-lived nature may still be sufficient to introduce stochastic resonance phenomena. 5 As we know from a set of different experiments addressing a variety of chemical reactions in cavities (see, e.g., Sec. 5.3), resonance effects (under normal incidence27) play a major role when modifying chemical properties under collective VSC. 22, 157, 158 Thus it comes as no surprise that periodic feedback effects between the frustrated (off-equilibrium) electronic-structure, the nuclear and cavity degrees of freedom and the thermal bath re-appear in a holistic theoretical description. 5 Considerable theoretical efforts went into investigating resonance effects with polaritonic reaction-rate theories 74, 159, 160, 161, 162, 163, 164, 165, 166, 167, 168 or few-molecule ab-initio simulations 169, 148, 170, but so far no consensus emerged with respect to the fundamental mechanism(s) at work. Clearly, the picture of spin glass-triggered resonance effects remains vague at the moment, and substantial theoretical and experimental effort will be required to unravel it and make it more quantitative and predictive. For the moment, we can only check if the implications of the SSK model agree qualitatively with experimental data.
@@
Figure 9: Connecting equilibrium and off-equilibrium picture for the SK (black) and SSK (red) spin glass assuming identical $q_{\rm EA}$. Left: Equilibrium order parameter distribution of a SK model (black) adapted from Ref. 171 in comparison with deviations of $q_{EA}(t)$ in a cavity. Right: In spin glasses two different (weak) off-equilibrium regimes (stationary and aging) can be distinguished for a small external field perturbation $h^{\prime}$, applied after a waiting time $t_{w}$. The stationary regime is governed by a linear fluctuation dissipation relation, which terminates at $C=q_{EA}$ given by the linear response susceptibility $S(q_{EA})=T\chi_{LR}$. In contrast, the aging regime is governed by the modified fluctuation dissipation relations in Eq. (55). It is bounded by the full thermal equilibrium susceptibility $S(0)=T\chi_{eq}$ at $C=0$. Notice that $\chi_{eq}\neq\chi_{LR}$ indicates replica symmetry breaking, whereas $\chi_{eq}=\chi_{LR}$ usually indicates normal hysteresis effects (red) that do not require the complex theory of spin glasses to describe its equilibrium properties. At first sight, the SSK model appears identical, i.e., trivial, in that picture. Nevertheless, its dynamics remains highly non-trivial, since it depends asymptotically on the waiting time according to Eq. (63). The light-blue region indicates that cavity-mediated correlations will deviate from an ideal SSK model, since the underlying probability distribution obviously depends on the chemical system and cannot be considered normally distributed. Corresponding investigation will require expensive numerical simulations.
@@
Figure 10: Experimental off-equilibrium measurement: The breakdown of fluctuation-dissipation relations (aging) measured the after cooling CdCr1.7In0.3S4 below the critical spin glass temperature of $T_{g}=16.2$ K. Bold symbols indicate the measured relaxation-correlation curve for different (finite) waiting times $t_{w}\in\{100,200,500,1000,2000\}$ [s]. Open symbols show the extrapolation to infinite waiting times $t_{w}\rightarrow\infty$The illustration was modified based on Fig. 7 Ref. 117, which contains experimental data from Ref. 143.
Figure 11: Illustrative sketch of classical nuclear trajectories evolving on a conservative potential energy surface (top), whereas aging effects (time-correlations) of the cavity-mediated correlation energy (bottom) may introduce non-conservative forces (orange arrows). The coupling to a thermal bath (Langevin equations), will result in canonical thermal equilibrium for the nuclei (top), whereas cavity-mediated spin glass effects are expected to give rise to (weak) non-canonical nuclear dynamics, prone to build up stochastic resonances (from Ref.5).
5.3 Experimental evidence
In the previous subsections we have seen that the emergence of a cavity-induced spin glass phase transition of the electronic-structure has several fascinating physical implications for an ensemble of molecules embedded in an optical cavity at ambient temperature $T$. In the following, we first compare the above theoretical predictions specifically with recent experimental nuclear magnetic resonance (NMR) results under VSC. 27 In Figs. 12 a-c the influence of VSC on the equilibrium concentration between two conformations of a molecular balance, sensitive to London dispersion forces, is studied with NMR spectroscopy. The experiments reveal that VSC can indeed modify the equilibrium rate constant and thus changes chemical properties that are directly related to electron-electron correlations, provided that the cavity is tuned on resonance with a specific C-H stretching mode of the solute molecules. If we disregard the resonance feature for the moment, the following three experimental observations seem to directly relate to our proposed spin/polarization-glass mechanism: First, the absence of cavity-induced chemical shifts indicates that the electronic-structure is on average not polarized by the cavity, which perfectly agrees with what we would anticipate from Eq. (17), i.e., the polarization glass. Second, the broadening of the chemical shifts seems modified under VSC, which one would expect for a cavity-induced polarization glass, i.e., from Eq. (18), or for modified rare events, i.e., for modified distribution tails that give rise to a spin glass phase. Notice, however, that in Ref. 27 the modified broadening is assigned to experimental artifacts and thus not further investigated. The third important insight of the NMR experiments is an abrupt change of the equilibrium constant at a specific collective strong coupling strength (influenced by the concentration) that does not scale further with the collective Rabi splitting. This suggests a phase transition at a specific collective coupling, as suggested by the SSK model. Notice further that while collective strong coupling (Rabi-splitting) has been reached in (dilute) gas phase experiments, so far no (local) change in chemistry was observed.172, 173, 174 This indicates that more condensed molecular ensembles (which possess significant dispersion effects) may be an crucial ingredient to change chemistry and thus modify matter-properties locally. This observation is inline with what we can derive from the quasi-dilute gas picture, which is an essential prerequisite for the formation of spin glass-like correlation effects of the electronic-structure.
In the following, we briefly look at a series of further experimental results that nicely connect, with what one would expect qualitatively from our mapping of the DSE electron correlations onto the SSK model.
  1. 1.
    Cavity-modified dispersion forces are known to play a crucial role in a series of different experiments, which report chemical changes under collective strong coupling. In particular, various cases have been reported where cavity-modified London-type dispersion forces27 alter the self-assembly of molecular structures under collective (and cooperative) VSC175, 176, 177. This macroscopic ordering across the molecular ensemble nicely illustrates the long-range nature of the all-to-all intermolecular DSE interactions.
  2. 2.
    Two different cases of phase-like transitions have been reported with respect to the collective coupling strength. For certain experiments (e.g., charge-transfer complexation178, supramolecular assembly of conjugated polymers175, conformational equilibrium constants27), an abrupt qualitative change is observed, with little dependency on the collective coupling strength beyond the critical point, whereas other cases seem to undergo a phase transition, with continuing parametric dependency (e.g. reaction rates22 or conductivity measurements in Refs. 179, 180). Notice, that typically the transition is not directly connected to the emergence of a Rabi splitting.158 Having in mind the abrupt change of the probability distribution of the free energy fluctuations of the SSK model, given in Eq. (60), different transition regimes must be anticipated. In particular, observables that rather depend on the mean of the distribution (i.e., bulk properties such as conformational equilibrium or self-assembly178, 175, 27) are expected to show only a weak parametric dependency. In contrast, rare events (e.g. chemical reactions,22 tunneling events180), that can rather be considered a measure for the tails of the probability distribution, are indeed sensitive to small changes of the collective coupling strength beyond the critical point. Remark: A phase-like transition on aggregation properties can also be observed for ground-state properties under electronic strong coupling. This suggests that our spin glass picture is transferable to electronic strong coupling situations.181 However, this aspect goes beyond the scope of this work.
  3. 3.
    Minimizing the DSE correlation energy (i.e., approaching the collective ground state) in Eq. (33) favors the delocalization of the intermolecular electronic orbitals. We anticipate that this effect should overall enhance tunneling effects180 and increase the conductivity182 or reactivity of the chemical ensemble under strong coupling along the polarization axis of the cavity. However, the impact of the DSE delocalization on chemical properties orthogonal to the polarization axis is difficult to anticipate. Depending on the chemical system, dispersion interactions may effectively be increased or suppressed when considering them spatially averaged. This could explain why it is delicate to control rare events, i.e., chemical reactions may either be suppressed22, 24 or increased.23 In that regard the symmetry properties of light and matter may play an important role and could help to make qualitative predictions.176
  4. 4.
    The inherent non-equilibrium nature of the SSK dynamics with no replica symmetry breaking suppresses many typical spin glass features, which makes the verification of our theoretical hypothesis particularly hard. Still, the trivially-modified fluctuation dissipation relations in Eq. (64) suggest an overall increase of the fluctuations (heating) in the aging regime and the emergence of cavity-modified hysteresis effects. Both effects have been reported in various experiments: For example, an effective temperature increase was measured for melting temperature of dsDNA,183, supramolecular polymerization175 or metal-to-insulator transitions in 1T-TaS2182 and for the large enhancement of ferromagnetism under VSC.184 Cavity-modified hysteresis is for example shown in Refs.184, 182. Notice, from a theoretical point of view, it may be surprising that an increase of the effective temperature can be accompanied by an ordering (self-assembly) of the system. However, as recently shown and explained generically in Ref.185, such an effect is characteristic for many complex system.
At this point it is important to note that the experimental evidence for our spin glass mapping remains vague and each of the mentioned aspect needs considerable research effort to be validated or falsified. In particular, the role of (stochastic) resonances and multiple time-scales remains unknown. Nevertheless, to our opinion many of the reported experimental results are in excellent qualitative agreement, with what one would anticipate from the physics of the SSK model.
Figure 12: Influence of VSC on the London dispersion-force-driven equilibrium, determined from NMR measurements according to Ref. 27: (a) Free energy difference between folded (1f) and unfolded (1u) conformer in benzened-$d_{6}$. (b) On- (red) and off-resonance (black) 1H NMR spectra of a 1 M solution. The shaded peaks allow to distinguish the two different conformers. The absence of a frequency shift between the two spectra indicates that on average the electronic-structure is not affected (no chemical shift). The magnitude of the peaks can be related to the equilibrium constant between the two conformers, which indicates VSC-induced chemical changes. (c) Varying the concentration (i.e., the collective coupling strength) reveals a critical concentration (dotted line), where suddenly a different conformational equilibrium constant is approached, which suggests the emergence of a cavity-induced phase transition.
Therefore, we believe that the spin glass-like nature of the dressed electronic-structure provides the most realistic and plausible theoretical framework that sets the necessary seed (instability mechanism) to trigger the resonance effects that have been observed in polaritonic experiments. Consequently, we consider the connection between polaritonic chemistry and the physics of a spin glass an excellent starting point to not only better understand current experimental findings, but also to stimulate novel theoretical and experimental directions. This should allow us to reach a holistic understanding of strongly-coupled light-matter systems. Moreover, our connection provides an interesting new direction for the field of spin glasses beyond condensed matter systems.
@@ 6 Summary and Conclusion
In this work, we highlight a novel and fundamental connection between two previously unrelated research domains: the emerging field of polaritonic chemistry and the well-established field of spin glasses and rare events in statistical mechanics. By mapping the cavity-mediated electron correlations in a quasi-dilute molecular ensemble under collective vibrational strong coupling (VSC) onto the spherical Sherrington–Kirkpatrick (SSK) model of a spin glass, our results suggest the emergence of a cavity-induced phase transition.
The theoretical pre-requisites for the proposed cavity-mediated spin glass phase can be summarized as follows:
@@
    Polarization ordering (a polarization-glass phase) emerges for sufficiently strong collective light-matter couplings to the cavity modes in the dilute gas limit. Resulting numerical convergence issues indicate a collectively-degenerate electronic ground-state for the molecular ensemble under VSC.
@@
    In the quasi-dilute gas limit, dipole self-energy(DSE)-modified electron-electron correlations $E_{\rm corr}^{\rm DSE}$ can be derived using the configuration interaction method, which dominate over the intermolecular Coulomb correlations. In that case, the intermolecular DSE-correlation energy has the form of a spin glass, provided that the density of states at the ground-state energy is sufficiently large. In particular, the DSE interactions are considered random due to the random orientation of the molecular ensemble with respect to the polarization of the cavity.
@@
    The DSE correlations connect to the SSK model when assuming $E_{\rm corr}^{\rm DSE}\sim H_{SSK}$. This suggests a spin glass-like phase transition induced by the cavity, provided that the fluctuations of the random DSE interaction are sufficiently strong. In particular, those fluctuations must be stronger than the thermal fluctuations of the dressed electronic correlations, i.e. $T<T_{c}$.
Entering a cavity-mediated spin glass phase using the SSK model suggests the following theoretical consequences for our polaritonic ensemble and polaritonic chemistry phenomena in general, which are in line with various experimental results (see Sec. 5.3):
@@
    The emergence of a spin glass-like phase transition, with respect to the collective light-matter coupling, confirms the relevance of cavity-mediated intermolecular dispersion effects. Moreover, the abrupt change of the local free-energy fluctuations occurs with respect to the number of collectively coupled degrees of freedom. This can be interpreted as a change in the probability distribution of ”local” dispersion effects, which suggests a qualitatively different behavior for bulk properties (e.g., self-assembly) and rare events (e.g., tunneling, reactivity) with respect to the collective coupling strength under VSC. Overall, the minimization of the DSE correlation energy favors the delocalization of the intermolecular orbitals along the polarization axis. However, perpendicular consequences (and symmetry in general176) must not be discarded to reach for a holistic chemical picture.
@@
    Thermal and non-adiabatic effects start to play a role for the collective electronic correlation problem in a cavity, even though these effects can usually be disregarded for an ensemble in free space. In particular, the SSK model suggests (extremely) long-lived time-correlations that effectively prevent the cavity-mediated electron correlations to reach equilibrium. As a consequence, an overall heating as well as changes of hysteresis effects can be anticipated from the SSK model, which is in line with experimental observations.
@@
    The non-equilibrium dynamics of the cavity-mediated electronic-structure implies non-conservative nuclear forces when assuming classical dynamics. Consequently, when coupling the classical degrees of freedom to a thermal bath (weakly-coupled Langevin dynamics), canonical equilibrium is no longer reached. Furthermore, non-conservative forces favor the build up of stochastic resonances. While a strong resonance behavior is known experimentally since the earliest experiments,22 considerable research effort will be required to disentangle the consequences of our spin glass picture on the formation of resonances under collective VSC.
Overall the SSK model suggests the existence of an instability of the dressed electronic subsystem, altering the established temporal and spatial scales to understand chemistry. Nevertheless, the interplay of the different scales (e.g. waiting time vs. stationary correlation vs. aging regime vs. inverse cavity frequency, …) still remains unexplored. The polarization spin glass picture can be considered as the limiting case of infinitely slow dynamics ($R,q_{\beta}$ ) (quasi-static picture). Overcoming this limitation will require further theoretical and computational efforts. In addition, determining chemically more realistic (heavily-tailed) probability distributions for the random interactions $J_{ij}$ will be challenging but important as well. The physical properties (e.g. scaling) of the resulting SSK model may deviate considerably from the case of a normally-distributed SSK model and it will be non-trivial to determine, even if realistic probability distributions are known. To gain a detailed understanding of all of the above aspects will require very interdisciplinary research efforts. However, we believe that pursuing this path will be extremely fruitful in many aspects. It may not only provide the missing theoretical piece to unravel the mysteries of polaritonic chemistry,131 but it could also trigger novel fundamental findings. For example, it could provide an theoretical and experimental tool to better understand and quantify rare events and nucleation (via the tails of $J_{ij}$). As we all know, chemistry, e.g., a chemical reaction, is an inherently rare off-equilibrium process, which is hard to describe apart from its rate. We therefore anticipate that cavity-controlled changes of the fluctuation-dissipation relations may eventually not only help to understand polaritonic chemistry, but they could provide novel insights into rare event and nucleation processes in general.
Biographies
Dominik Sidler, born in Ebikon, Switzerland, is a group leader at the Max-Planck Institute for the Structure and Dynamics of Matter (MPSD) in Hamburg, Germany. He is also a Scientist at Paul Scherrer Institute (PSI) in Villigen, Switzerland. He obtained his Ph.D. in computational chemistry from ETH Zürich, Switzerland, in 2018. Afterwards, he moved as a postdoctoral fellow to the MPSD in Hamburg, Germany. In 2024, he returned as a Scientist to PSI Switzerland and was appointed research group leader at MPSD. His research group works on the theoretical and computational description of (strong) light-matter interactions in polaritonic chemistry. He is a team member of the international ERC Synergy project ”UnMySt” and his current research focus is on first-principles approaches to describe emergent local and collective strong coupling phenomena in optical cavities, as well as the impact of vacuum field fluctuations on molecular non-equilibrium conditions at finite temperatures.
Michael Ruggenthaler is research group leader at the Max-Planck Institute for the Structure and Dynamics of Matter in Hamburg, Germany. He obtained his Ph.D. in theoretical physics jointly from the Max-Planck Institute for Nuclear Physics and the University of Heidelberg, Germany, in 2009. As Schrödinger Fellow of the Austrian Science Fund he worked at the University of Jyväskylä, Finland, and then established an independent research group at the University of Innsbruck, Austria, before he moved to the Hamburg in 2016. His research group works on the theoretical and mathematical foundations of quantum many-body theories, in and out of equilibrium. He has been involved in several successful international research collaborations such as the QuantEra project ”RouTe” or the ERC Synergy project ”UnMySt”. His current research focus is on first-principles approaches to ab initio quantum electrodynamics, polaritonic chemistry and cavity-modified material properties.
Angel Rubio, born in Oviedo, Spain, is the Director of the Theory Department of the Max Planck Institute for the Structure and Dynamics of Matter. He is also Distinguish Research Scientist at the Simons Foundation’s Flatiron Institute (NY, USA) and Professor/Chair for condensed matter physics at the University of the Basque Country in Donostia-San Sebastián, Spain. He received his PhD in Physics with honors from the University of Valladolid (UVA) in 1991 where he did fundamental work on the structural and optical properties of metallic clusters. Then moved to a postdoctoral position at UC Berkeley-Physics (92-95) where he predicted a new type of boron-nitride nanotubes triggering their ensuing experimental synthesis. Between 1994 and 2001 as Professor at UVA he started the ab initio materials research open-source project octopus used now by over 1000 groups worldwide. Diverse Professorships at École Polytechnique Paris, FU Berlin and Montpellier followed. In 2001 he moved as Chair of Condensed Matter Physics at UPV/EHU in San Sebastian. There he engaged in highly successful work on modeling of excited-state properties of materials and nanostructures setting the foundations of modern theoretical spectroscopy. In August 2014 he accepted the position as Max Planck Director. There he has pioneered the development of quantum electrodynamical density functional theory (QEDFT), a novel theoretical framework for strong light-matter phenomena in chemistry and materials sciences . His work has been recognized by several awards, including the 2023 Spanish National Physics Prize “Blas Cabrera” 2018 Max Born medal and prize, 2016 Medal of the Spanish Royal Physical Society and the 2014 Premio Rey Jaime I for basic research, and more, and elected member of different academies, including the German Leopoldina Academy and Berlin-Brandenburgischen Akademie der Wissenschaften, the European Academy of Sciences, the Academia Europaea, and a foreign associate member of the National Academy of Sciences (USA).
@@
We gratefully acknowledge all members of our ERC synergy grant proposal team (Tal Schwartz, Thomas Ebbesen, Abraham Nitzan, Sharly Fleischer, Cyriaque Genet and Maxim Sukharev) for inspiring discussions and helpful comments. Furthermore, we particularly thank Benjamin Lev, Markus Mueller, Dries Sels and Antoine Georges for their valuable and critical feedback on the spin glass aspects. Finally, we would like to thank all four referees for their very constructive feedback. Specifically, the comments of referee 3 were extremely helpful to connect our predictions to experimental observations. This work was made possible through the support of the European Research Council (ERC-2024-SYG-101167294, UnMySt), the Cluster of Excellence Advanced Imaging of Matter (AIM), Grupos Consolidados (IT1249-19) and SFB925. We acknowledge support from the Max Planck-New York City Center for Non-Equilibrium Quantum Phenomena. The Flatiron Institute is a division of the Simons Foundation.